### v1

# AI-Based Detection of In-Treatment Changes from Prostate MR-Linac Images

Abstract
Purpose: To investigate whether routinely acquired longitudinal MR-Linac images can be leveraged to characterize treatment-induced changes during radiotherapy, particularly subtle inter-fraction changes over short intervals (average of 2 days).
Materials and Methods: This retrospective study included a series of 0.35T MR-Linac images from 761 patients. An artificial intelligence (deep learning) model was used to characterize treatment-induced changes by predicting the temporal order of paired images. The model was first trained with the images from the first and the last fractions ($F_{1}$-$F_{L}$), then with all pairs (All-pairs). Model performance was assessed using quantitative metrics (accuracy and AUC), compared to a radiologist’s performance, and qualitative analyses—the saliency map evaluation to investigate affected anatomical regions. Input ablation experiments were performed to identify the anatomical regions altered by radiotherapy. The radiologist conducted an additional task on partial images reconstructed by saliency map regions, reporting observations as well. Quantitative image analysis was conducted to investigate the results from the model and the radiologist.
Results: The $F_{1}$-$F_{L}$ model yielded near-perfect performance (AUC of 0.99), significantly outperforming the radiologist. The All-pairs model yielded an AUC of 0.97. This performance reflects therapy-induced changes, supported by the performance correlation to fraction intervals, ablation tests and expert’s interpretation. Primary regions driving the predictions were prostate, bladder, and pubic symphysis.
Conclusion: The model accurately predicts temporal order of MR-Linac fractions and detects radiation-induced changes over one or a few days, including prostate and adjacent organ alterations confirmed by experts. This underscores MR-Linac’s potential for advanced image analysis beyond image guidance.
Abbreviations
EBRT = external beam radiotherapy, RT = radiation therapy, MRI = magnetic resonance imaging, MR-Linac = Magnetic Resonance-Linear Accelerator, ADC = apparent diffusion coefficient, AI = artificial intelligence, AUC = area under the receiver operation characteristic curve, CI = confidence interval, PSA = prostate-specific antigen, ADT = androgen deprivation therapy
Summary
MR images from MR-Linac fractions contain therapy-induced changes that can be systematically detected using artificial intelligence.
Key Points
  1. 1.
    An artificial intelligence model accurately determined the temporal order of images from MR-Linac fractions, exceeding human expert performance and demonstrating the presence of radiation therapy-related changes.
  2. 2.
    Radiation therapy–induced changes were most pronounced in the prostate, with additional contributions from the bladder and surrounding pelvic tissues, supporting the clinical value of MR-Linac imaging for treatment monitoring.
Keywords
Prostate Cancer, MR-Linac, Radiation Therapy, Longitudinal Image Analysis, Adaptive Radiotherapy, Deep Learning, Artificial Intelligence
1 Introduction
A common therapeutic strategy for prostate cancer patients is external beam radiotherapy (EBRT), delivered either independently or in combination with surgery or hormone therapy [18, 17]. Image-guided radiation therapy (RT) has become standard practice for ensuring precise localization of the target and safe treatment delivery. Magnetic Resonance-Linear Accelerator (MR-Linac) systems combine MRI with a linear accelerator, enabling the acquisition of high–soft-tissue–contrast MR images at every treatment fraction [22, 20]. This integration allows visualization of the prostate, bladder, rectum, and adjacent tissues before radiation delivery and day-to-day anatomical changes throughout the treatment course [19, 22, 20]. Moreover, the MR images allow for adaptive therapy taking inter- and intra-fraction motion management into account, modifying each treatment fraction as needed based on spatiotemporal variations to ensure maximal target dose delivery with minimal dose to adjacent organs [15, 25, 22, 7].
While MR-Linac provides unprecedented access to longitudinal MR imaging throughout the treatment course, potential treatment-induced changes in the images have been unexplored. This is due to the relatively poor image quality and the subtle changes that are challenging for human observers to detect, combined with the fact that these images are primarily used for guiding RT and rarely reviewed beyond gross anatomical assessment. As a result, despite their potential clinical relevance for adaptive planning, treatment personalization, and toxicity prediction, systematic characterization of fraction-to-fraction changes remains limited [7].
Several studies have reported significant prostate volume changes in diagnostic-quality MR images during a course of RT [17, 1]. Longitudinal studies using multiparametric MRI [2, 9] and quantitative MRI [27] have demonstrated changes in prostate tumors and/or surrounding tissues before and after treatment delivery. From MR-Linac images, prior works have shown changes in apparent diffusion coefficient (ADC) [3, 8]. These prior studies suggest that MR imaging during RT can capture dynamic, patient-specific morphological and tissue changes. However, to the best of our knowledge, inter-fraction changes in routinely acquired MR-Linac RT imaging have not been previously studied.
In this study, we developed an AI (artificial intelligence) model designed to detect subtle changes in MR images from MR-Linac fractions in one or a few days, adapting a longitudinal image comparison framework [12]. We hypothesized that a model capable of accurately ordering images chronologically would rely on features that evolve consistently with cumulative radiation exposure. The model achieved over $95\%$ accuracy, substantially outperforming an expert, and identified prostate and surrounding anatomical structures, such as bladder and pubic symphysis, as key regions driving radiation-induced changes. These findings are supported by input ablation experiments and quantitative image analysis. Overall, our study suggests that beyond its established role in patient alignment and dose planning, MR-Linac imaging can be leveraged to study longitudinal, treatment-induced changes during RT, providing AI-derived imaging features that may support treatment monitoring and adaptation.
2 Materials and Methods
2.1 Data
The study is HIPAA-compliant and IRB-approved with a waiver of informed consent for retrospective studies. All data were de-identified prior to analysis.
We retrospectively collected $1,071$ patients who underwent MR-guided RT from 2018 to 2025, screened to 761 patients for the study (Details of selection are in the Supplementary Section A1 and flow diagram is in Supplementary Figure A1). The images include simulation scans (Sim in Figure 2A) and corresponding treatment fractions ($F_{1}$ to $F_{n}$ in Figure 2A)), with inter-fraction intervals of one day or more (Supplementary Figure A2). All images were acquired using a ViewRay MRIdian MR-Linac system (ViewRay Inc., USA) equipped with a 0.35-T MRI, using True Fast Imaging with Steady-State Free Precession (TRUFI) sequence. The in-plane pixel spacing was $1.5\times 1.5~\mathrm{mm}^{2}$, with a slice thickness of $1.5$ or $3~\mathrm{mm}$. Apart from cropping centered on prostate masks (see Supplementary Section A2 for segmentation method) to (80, 80, 80) sized volumes, no further pre-processing was performed.
Data were randomly split on a patient-wise basis into training (60%, n = 457), validation (20%, n = 152), and test data (20%, n = 152). All MR images were cropped to dimensions of (80, 80, 80), centered at the centroid of the prostate mask (see Supplementary Section  A2 for segmentation method).
2.2 Model
The overall framework (Figure 2B) is based on learning-based Inference of Longitudinal imAge Changes (LILAC) [12], which learns meaningful changes from paired longitudinal images via predicting temporal order. As temporal ordering has demonstrated effectiveness in capturing monotonic changes, we adopted it to identify features consistently changing over the course of treatment. We employed a Siamese 3D convolutional neural network [5, 6, 13] with ResNet-18 [10] architecture. The model outputs logits for binary classification, predicting whether the input pair is in the correct or reversed temporal order. These logits reflect the model’s confidence, informed by learned differences between paired longitudinal images. Code is provided in https://github.com/anonymized/mr-linac-ai/will_be_added. Further details are available in Supplementary Section A3.
Figure 1: Overview of the MR-Linac workflow and AI framework.
Figure 2: (Continued) (A) Process of MR-Linac sessions across multiple fractions. Sim refers to the simulation scans obtained prior to MR-Linac treatment sessions. $F_{1}$, $F_{2}$, …, $F_{L}$ represent MR images acquired at the first, second, …, and last fractions, respectively, prior to image-guided radiation therapy (RT). (B) Schematic diagram of the LILAC-based AI model. Paired 3D MR image volumes from fractions are input to a Siamese 3D convolutional neural network with ResNet-18 [10] architecture, in which features of each image are obtained. The differences between these two features are fed into a fully connected layer. The logits are used for binary classification to predict if the image pairs are in the correct temporal order or not. (C) Experimental design for evaluation. A radiologist performed the same temporal-ordering task first with full volumes and next with saliency-restricted images. The radiologist also evaluated anatomical regions highlighted by the saliency maps. Model input ablation experiments were designed to evaluate altered anatomical regions. Performance changes when the prostate, bladder, or both regions or the surrounding box regions were masked in MR images (Organ Masked, Box Masked) were evaluated, as well as performance for partial MR images encompassing the organ regions (Only Organ). Performance for using only masks (Mask) was evaluated to investigate the effects of the organ shapes, compared to MR intensities.
For training, MR images from multiple fractions were paired in all possible ordered permutations. A curriculum learning was adopted, where initial training used only first-last fraction image pairs ($F_{1}$$F_{L}$ pairs), which exhibit the largest radiation-induced differences and are therefore easier to learn [4]. The model was further trained using all combinations of image pairs (All pairs; All-pairs model), with parameters initialized from those of the best $F_{1}$$F_{L}$ model. More details are in Supplementary Section A4.
2.3 Assessment of Changes Across Fractions via Model Performance
The held-out test data ($152$ patients, $732$ image pairs) were used to evaluate the performance across all experiments.
Temporal ordering performance evaluation.—
Model performance was assessed using accuracy and the area under the receiver operating characteristic curve (AUC), with confidence intervals (CI) obtained from $1,000$ bootstrap resamples of the test data.
Comparison to expert: A radiologist with over 15 years of experience performed pairwise image comparison task to identify the correct order for the same $F_{1}$-$F_{L}$ pairs.
Sim-F1 test: As $Sim$ and $F_{1}$ images are acquired before treatment, these pairs (SimSim-$F_{1}$) were used to evaluate the model’s sensitivity to radiation-induced changes rather than temporal changes alone.
Fraction-pairwise performance evaluation.—
As shown in Kim et al. [13], the model output reflects the magnitude of changes between image pairs. Based on this observation, we examined logits of the All-pairs model across fraction levels. To further quantify the relationship between fraction intervals and logits, we computed the Pearson correlation coefficient and fit a linear mixed-effects model with patient as a random effect (See Supplementary Section A5 for more details).
2.4 Identification of Anatomical Regions with Changes Across Fractions Using Model
Saliency map.—
Saliency maps highlight regions that the model identifies as contributing most strongly to the differences between a given image pair, which are likely to reflect RT-induced effects. We used a modified Gradient-weighted Class Activation Mapping (Grad-CAM) [23, 12] to calculate the saliency maps. Implementation details and population-level visualization on an atlas are available in Supplementary Section A6.
Input ablation.—
Input ablation experiments were conducted to identify regions critical for the model’s prediction by examining model performance changes depending on input changes (Figure 2). Occlusion was performed using either the segmented regions (see Supplementary Section A2 for segmentation method) themselves (organ-masked-MR input) or bounding boxes surrounding the segmented areas (box-masked-MR input), where lower performance indicates stronger model reliance on the occluded organ regions. Preservation was performed by retaining only the prostate, bladder, or both segmented regions in the MR images (only-organ-MR input), where higher performance indicates stronger model reliance on the preserved organ regions. These ablated inputs retain both organ size and intensity. To further assess the decoupled effect of organ size from intensity, segmented organ masks (prostate, bladder, or both) were used (mask input).
Characteristic Training Data Validation Data Test Data
No. of patients 457 152 152
Age (y) 73 (10) 73 (11) 72.5 (10.25)
No. of fractions per patient 5 (0) 5 (0) 5 (0)
Days between fractions 2 (1) 2 (1) 2 (1)
PSA level (ng/mL) 2.36 (4.38) 2.13 (4.02) 2.21 (2.59)
No. of patients with ADT 180 (39) 66 (43) 55 (36)
Note.—Age, the number of fractions per patient, days between fractions, and PSA level are reported as medians with interquartile ranges in parentheses. The number of patients with ADT is reported as medians with percentages in parentheses. No significant differences were found among values from the training, validation, and test data, except for the number of patients (independent two-sided t-test).
Table 1: Patient Characteristics.
Pair Performer ACC AUC
$F_{1}$-$F_{L}$ AI 0.95 (0.93, 0.98) 0.99 (0.99, 1.00)
Expert 0.82 (0.75, 0.87) -
All-pairs AI 0.91 (0.90, 0.92) 0.97 (0.96, 0.97)
AI, inference for $Sim$-$F_{1}$ pairs 0.39 (0.34, 0.45) 0.34 (0.28, 0.40)
Note.— Significantly different from the bolded results (independent two-sided t-test, $p<.0001$)
Table 2: Performance of the model and the expert.
Expert Saliency-Restricted MR Evaluation
The radiologist performed the same temporal-order identification task using images restricted to the regions highlighted by the saliency maps, with the saliency maps themselves not shown (saliency-restricted-MR; implementation details in Supplementary Section A7). The radiologist also documented the rationale for each decision. The overlaid saliency maps were later provided for expert’s evaluation in order to interpret the model’s decision-making process and explain regions affected by RT.
2.5 Image Analysis
Quantitative image analysis was performed to further investigate RT-induced changes and to compare them with the model and expert evaluations. Changes in organ volume, and in the mean and standard deviation of image intensity for the prostate and bladder were assessed for all patients between $F_{1}$ and $F_{L}$. Prostate and bladder were defined using inferred masks generated by the nnU-Net model (Supplementary Section A2).
2.6 Statistical Analysis
An independent two-sided t-test was conducted for comparisons of scores (accuracy and AUC) and patient characteristics. The Wilcoxon signed-rank test was used to compare organ volume and intensity from $F_{1}$ to $F_{L}$. For all statistical tests, p-value of 0.05 was used as the threshold for statistical significance. See Supplementary Section A8 for details on the tools used.
3 Results
Figure 3: Pairwise performance (A) Model logits of the All-pairs model for different image pairs. Model logits reflect the model’s confidence in its predictions, which can be interpreted as the magnitude of changes detected between paired images. Pairs of numbers on the x-axis indicate the fractions at which the paired images were acquired. Cases were grouped by the first fraction and ordered and color-coded by the interval between the paired fractions. Stars indicate statistical significance between compared cases (independent two-sided t-test; ns: .05<p.05<p, **: $01<p\leq.01, ***: .00$$001<p\leq.001$, ****: $p\leq.0001$). Means within the same first fraction group are shown in red, whereas means with the same fraction interval but different first fraction groups are shown in green. (B) Linear mixed-effects model showing the relationship between fraction intervals from the first fraction and model logits (black), with patient-wise gradients shown in multicolor.
Index Expert Evaluation of Grad-CAM Peak # of Cases
1 Bladder lumen 31 (20.53%)
2 Pubic symphysis 29 (19.21%)
3 Bladder lumen, public symphysis, and periprostatic fascia (1 + 2 + 8) 18 (11.92%)
4 Pubic symphysis and bladder lumen (1 + 2) 17 (11.26%)
5 Base of prostate at urethral inlet 8 (5.30%)
6 Public symphysis and base of prostate at urethral inet (2 + 5) 8 (5.30%)
7 Pubic symphysis and periprostatic fascia (2 + 8) 7 (4.64%)
8 Periprostatic fascia 6 (3.97%)
9 Bladder wall 5 (3.31%)
10 Bladder lumen and lateral prostate (1 + 14) 4 (2.65%)
11 Random fascia above the bladder 4 (2.65%)
12 Urethral inlet and periprostatic fascia (5 + 8) 3 (1.99%)
13 Bladder apex 2 (1.32%)
Note.—Regions with only one case are omitted from the table: prostate peripheral zone, pubic symphisis and lateral prostate possibly peripheral zone (2 + 14), anterior prostate and possibly artifact, pubic symphysis and tissue above bladder (2 + other), urethral inlet and bladder lumen (1 + 5), like 3 but ill-defined, ill-defined areas spanning the pubic symphysis and prostate base with additional bilateral areas in fascia above the bladder, periprostatic fascia and fascia above the bladder (8 + 11), and bladder lumen and periprostatic fascia (1 + 8).
Table 3: Expert evaluation of peak regions of saliency maps from the All-pairs model on $F_{1}$-$F_{L}$ pairs.
3.1 Patient Characteristics
All patients in the cohort (Table 1) were diagnosed with elevated prostate-specific antigen (PSA) levels and prostate cancer. All patients except 43 patients received five fractions of MR-Linac treatment, mostly administered every two days (Supplementary Figure A2). Approximately 40% of patients in the cohort underwent androgen deprivation therapy (ADT). There were no statistically significant differences between the train, validation, and test splits.
3.2 Evaluation of Changes in Images across Fractions via Model Performance
Because temporal order is inferred from image differences, output logits were used as a surrogate measure of change across RT fractions.
Model performance and comparative analysis.—
The model performance on the test set was quantitatively evaluated (Table 2). The $F_{1}$-$F_{L}$ model yielded the accuracy of 0.95 (95% CI: 0.93, 0.98), and the AUC of 0.99 (95% CI: 0.99, 1.00). The model outperformed the expert achieving an accuracy of 0.82 (95% CI: 0.76, 0.88), of which the difference was statistically significant (independent two-sided t-test, $p<.0001$).
The All-pairs model achieved an accuracy of 0.95 (95% CI: 0.93, 0.97) and an AUC of 0.99 (95% CI: 0.99, 1.00). In contrast, inference on $Sim$-$F_{1}$ pairs yielded a significantly lower accuracy of 0.40 (95% CI: 0.34, 0.45) and AUC of 0.34 (95% CI: 0.28, 0.40) (independent two-sided t-test, $p<.0001$). The model achieved high accuracy on images acquired during RT but performed poorly on pre-RT images (SimSim-$F_{1}$), indicating that RT-driven changes underlie the model’s decisions. Patients who underwent ADT showed no statistically significant differences in model logits (independent two-sided t-test), further supporting that predictions are driven by RT-induced changes.
Figure 4: Saliency map on an atlas. The atlas was built from $F_{1}$ of all patients in the test data. The visualized heatmap is the mean of saliency maps for all patients in the test data, obtained from All-pairs model inference on $F_{1}$-$F_{L}$ pairs and transformed into the atlas space. Column titles indicate the slice orientations (axial, sagittal) and the primary regions of interest highlighted by the saliency maps, although these regions span multiple organs.
Model performance on detecting RT-induced changes.—
Image differences between RT fractions are expected to increase with longer fraction intervals due to potential cumulative radiation exposure. To examine this, the performance of the All-pairs model for each fraction pair was evaluated (Figure 3 and Supplementary Figure A3). Model logits were positively correlated with the inter-fraction interval (Pearson $r=0.59$, $p<.0001$), as well as accuracy and AUC all increased proportionally with the fraction intervals (Supplementary Figure A3). Within groups of image pairs sharing the same first image (e.g. 1-2, 1-3, 1-4, and 1-5), the mean model logits increased with the fraction interval (red lines in Figure 3A). Among adjacent image pairs (1-2, 2-3, 3-4, and 4-5), the $F_{2}$-$F_{3}$ pairs exhibited significantly higher than other one-interval pairs (green lines in Figure 3C), suggesting that RT-induced changes are most pronounced between the second and third fractions. These results demonstrate that the model predictions can quantify RT-induced changes between fractions.
Incorporating individual variability in the linear mixed-effects model resulted in a significant difference from the reduced model without random effects, according to the likelihood ratio test ($p<.0001$) (Figure 3). This indicates that the changes vary across patients, suggesting heterogeneous treatment effects.
3.3 Localization of RT-induced changes driving model predictions
Saliency Map.—
Modified Grad-CAM[23, 13] based saliency maps were used to interpret which regions contribute to the model’s predictions. According to the radiologist’s evaluation, the saliency peaks most frequently occurred in the bladder lumen (20.53%) and pubic symphysis (19.21%), with the prostate commonly involved (Table 3). Group-level visualization of the heatmaps on the atlas of all patients is shown in Figure 4 (Peak distribution of the heatmaps in Supplementary Figure A4), of which the primary peak regions are the pubic symphysis and bladder, while the area spans multiple anatomical regions, including the prostate.
Input Ablation.—
Input ablation experiments were conducted to identify anatomical regions critical to the model’s performance (Table 4). All input ablation test results were significantly different from the baseline result. Occluding both prostate and bladder regions in MR (Organ-masked-MR and Box-masked-MR) shows a significant drop in the performance, indicating that prostate, bladder, and surrounding box regions contain important RT-related changes. Box-masked-MR caused greater performance degradation than Organ-masked-MR, possibly suggesting the significance of shape changes in model prediction, which motivated the evaluation of Mask input.
Preservation experiments show the effect of organ shape and texture as input (Only-organ-MR) and only shape as input (Mask). Using masks alone (Mask) reduced performance relative to Only-organ-MR, although accuracy remained above 0.9 ($F_{1}$$F_{L}$ model) and 0.8 (All-pairs model). Prostate mask had a larger impact than bladder mask. The results suggest that prostate size is a major contributor, but intensity also contributes. The performance degradation relative to the baseline was greater for All-pairs model than $F_{1}$$F_{L}$ model, indicating that shape and texture information are essential to capture subtle inter-fraction changes.
All-pairs model performed best when both prostate and bladder were preserved, whereas $F_{1}$$F_{L}$ model achieved higher performance when only prostate was preserved. This indicates that the most pronounced changes between pre- and post-treatment are localized to the prostate, whereas more subtle inter-fraction changes involve both prostate and bladder.
$F_{1}$-$F_{L}$ model All-pairs model
Input Type Organ ACC AUC ACC AUC
Baseline MR - 0.95 $\pm$ 0.01 0.99 $\pm$ 0.00 0.91 $\pm$ 0.01 0.97 $\pm$ 0.00
Organ-masked-MR Prostate -0.03 $\pm$ 0.01 -0.01 $\pm$ 0.00 -0.01 $\pm$ 0.01 -0.00 $\pm$ 0.00
$\downarrow$ Bladder -0.02 $\pm$ 0.01 -0.01 $\pm$ 0.00 -0.04 $\pm$ 0.01 -0.02 $\pm$ 0.00
Prostate $\cup$ Bladder -0.05 $\pm$ 0.02 -0.04 $\pm$ 0.01 -0.04 $\pm$ 0.01 -0.03 $\pm$ 0.00
Box-masked-MR Prostate -0.07 $\pm$ 0.02 -0.06 $\pm$ 0.02 -0.04 $\pm$ 0.01 -0.02 $\pm$ 0.00
$\downarrow$ Bladder -0.04 $\pm$ 0.02 -0.02 $\pm$ 0.01 -0.08 $\pm$ 0.01 -0.06 $\pm$ 0.01
Prostate $\cup$ Bladder -0.20 $\pm$ 0.02 -0.19 $\pm$ 0.02 -0.25 $\pm$ 0.01 -0.25 $\pm$ 0.01
Only-organ-MR Prostate -0.03 $\pm$ 0.02 -0.01 $\pm$ 0.01 -0.10 $\pm$ 0.01 -0.08 $\pm$ 0.01
$\uparrow$ Bladder -0.09 $\pm$ 0.02 -0.06 $\pm$ 0.01 -0.14 $\pm$ 0.01 -0.12 $\pm$ 0.01
Prostate $\cup$ Bladder -0.03 $\pm$ 0.02 -0.02 $\pm$ 0.01 -0.08 $\pm$ 0.01 -0.05 $\pm$ 0.00
Mask Prostate -0.05 $\pm$ 0.02 -0.04 $\pm$ 0.01 -0.16 $\pm$ 0.01 -0.15 $\pm$ 0.01
$\uparrow$ Bladder -0.21 $\pm$ 0.03 -0.22 $\pm$ 0.03 -0.29 $\pm$ 0.01 -0.30 $\pm$ 0.01
Prostate $\cup$ Bladder -0.09 $\pm$ 0.02 -0.07 $\pm$ 0.02 -0.12 $\pm$ 0.01 -0.13 $\pm$ 0.01
Note.—Values in the first row are presented as mean $\pm$ standard deviation. All other values represent the difference from the baseline mean $\pm$ standard deviation.
: All cases are significantly different from the baseline results (independent two-sided t-test, $p<.0001$).
$\downarrow$: Performance degradation compared to the baseline indicates strong model dependence on the targeted organ regions. The worst case is highlighted in bold.
$\uparrow$: Higher performance indicates strong model dependence on the targeted organ regions. The best case is highlighted in bold.
Table 4: Input ablation.
Expert Saliency-Restricted-MR Evaluation.—
The expert additionally assessed the model’s saliency maps through performing a prediction test on saliency-restricted MR images (Prediction Test on Saliency-Restricted-MR in Figure 2C). Although the FOV of the images is limited, the expert achieved an accuracy of 0.72 (95% CI: 0.66–0.79), showing relevance of model-highlighted regions for capturing changes.
For full volume prediction test, the expert specified the decision criteria in descending order of importance: (1) darkening of the prostate, (2) reduced distinction between the peripheral zone and transition zone borders, (3) enlargement of the prostate, and (4) reduced definition of the prostate’s outer edge. The expert primarily relied on changes in prostate volume and MR image characteristics, including prostate brightness and clarity of the prostate edges. The rationales were similar for saliency-restricted-MR test (Supplementary Table A1).
3.4 Organ Volume and Intensity Changes in Images During RT
Changes in prostate and bladder volume and intensity between $F_{1}$ and $F_{L}$ were analyzed to validate the model and expert findings (Figure 5). Prostate volume increased significantly from $F_{1}$ to $F_{L}$, whereas bladder volume decreased significantly (one-sided Welch’s t-test, $p<.0001$). Both organs became darker, with the mean intensity in each region decreasing significantly from $F_{1}$ to $F_{L}$ (independent one-sided t-test, $p=.02$ and .001.001, respectively). Additionally, the variance of intensity, reflecting heterogeneity, increased significantly in the prostate, whereas it decreased significantly in the bladder (one-sided independent t-test, $p=.002$ and $p<.0001$, respectively).
The observed increase in prostate volume and reduction in brightness aligns with both the model’s results and the expert’s assessments. These results suggest that the model’s detection of inter-fractional changes may be partially explained by changes in prostate volume and intensity.
4 Discussion
Figure 5: Changes in volume and intensity of prostate and bladder in MR images. An increase in the difference between the two means is indicated in red, whereas a decrease is indicated in blue. Stars indicate statistical significance of the difference (Wilcoxon signed-rank test; ****: $p\leq.0001$)
In this study, we show that AI–based analysis of prostate MR-Linac images, which are primarily used to guide radiotherapy, has potential clinical value for detecting in-treatment anatomical and tissue changes.
This study characterizes therapy-induced changes in MR-Linac images across fractions over short time scales—one or a few days between consecutive fractions. Analyzing changes in MR-Linac images is challenging and has received limited attention, as many inter-fraction changes are subtle and difficult to detect visually. We not only developed an AI model that accurately predicts the temporal order of MR images acquired during MR-Linac treatment fractions separated by one to several days apart, but also conducted extensive evaluations—including model input ablation and clinician assessments—to investigate in-treatment changes and validate the model’s clinical relevance.
The model achieved near-perfect performance, surpassing expert assessment, suggesting the presence of subtle differences that are challenging to detect visually (Table 2). Performance increased with inter-fraction interval, consistent with cumulative radiation effects (Figure 3). No significant performance difference was observed between patients with and without ADT, suggesting that RT was the primary driver of the model’s decisions. Failure cases occurred primarily in consecutive fraction pairs and did not correspond to any identifiable patient characteristics (e.g., ADT) or imaging parameters (e.g., slice thickness).
Anatomical regions detected by the model were investigated through saliency maps and input ablation. The saliency maps mainly highlighted the pubic symphysis, bladder, and prostate (Table 3, Figure 4). This is consistent with reports that some patients who had radiotherapy or MR-Linac treatment for prostate cancer experience genitourinary symptoms such as urinary frequency, urethral stricture, and cystitis [16, 21, 28]. While less has been reported on the pubic symphysis, pubic bone osteomyelitis and changes in T2- and T1-weighted signals of the involved pubic rami were reported in patients who had RT for prostate cancer [24]. The result suggests that the AI model highlights RT-affected regions, including unexpected side effects, supporting its potential use in treatment assessment.
Input ablation revealed that the prostate is the primary contributor to the prediction and the bladder and surrounding regions (bounding boxes) of both organs also contribute, indicating that these areas are affected by RT. While organ size is the dominant factor, MR intensity also plays a role. The model primarily leveraged changes in the prostate and bladder for inter‑fraction predictions, with pre‑ to post‑treatment differences most pronounced in the prostate.
Quantitative analysis confirmed that the prostate enlarged and darkened with increased intensity heterogeneity, while the bladder shrank, darkened, and intensity became more uniform (Figure 5). RT may have contributed to prostate enlargement and associated bladder compression, possibly reflecting treatment-related inflammatory changes. Signal darkening occurred in both, but increased intensity variability was observed only in the prostate, suggesting differential tissue responses to targeted versus non-targeted radiation.
When the expert conducted the same test on saliency-restricted-MR images, the expert yielded performance significantly different from the baseline result but above chance, indicating that the model’s saliency maps capture some information about inter-fractional changes. The expert primarily relied on prostate features—volume, darkening, and edge clarity (Table A1), whereas the model also emphasized regions around the bladder and pubic symphysis (Table 3, Figure 4), suggesting it may detect additional subtle changes not readily observed by humans.
Several limitations should be acknowledged. Grad-CAM may yield broad activation patterns; fine-grained explainability methods could improve localization. This proof-of-concept used data from a single platform, and performance on other platforms remains to be determined.
In conclusion, this investigation provides conclusive evidence that AI can accurately determine the temporal order of treatment scans, suggesting the feasibility of AI-based detection of in-treatment changes as a potential biomarker of treatment response or adverse effects. Future work will serve to evaluate the ability of the model to predict biochemical failure and the development of adverse effects with the goal of appropriately selecting patients for primary and adjuvant therapy.
Supplementary Material
A1 Patient screening
Of the 1,071 patients, those who underwent multiple treatment sets (n = 57) and those without a prostate or prostate mask (n = 253) were excluded (Supplementary Figure A1), as their imaging characteristics may differ from the rest of the cohort.
A2 Segmentation of Organs in MR Images from MR-Linac Fractions
The prostate, bladder, and rectum are routinely contoured in RT guide images obtained before treatment by radiation oncologists for treatment planning purposes. Since contouring is routinely performed only on the guide images, we trained an nnU-Net [11] on the guide images to get prostate, bladder, and rectum masks in MR images from MR-Linac fractions. We assumed that the model would generalize to images from fractions, as both were obtained from the same MR-Linac system and target the same anatomical regions with consistent imaging protocols.
The segmentation model was trained, validated, and tested using split datasets of guide MR images with corresponding prostate, bladder, and rectum masks delineated by radiation oncologists. On the test set, the Dice scores of the model were 0.85, 0.95, and 0.82 for prostate, bladder, and rectum. The trained model was then applied to MR images acquired during MR-Linac treatment fractions to generate organ masks. Partial qualitative evaluation of the resulting masks, together with quantitative comparison between the trained MR-Linac temporal-ordering models, demonstrated that this approach outperformed masks obtained via image registration.
We used dilated masks for input ablation experiments, except for the mask-only condition, to ensure that organ boundaries were included in the occluded or preserved regions. Binary masks were morphologically dilated by two voxels in all directions using a 3D 26-connected structuring element.
A3 Model Architecture Details
The overall framework (Figure 2B) is based on learning-based inference of longitudinal image changes (LILAC) [12], which extracts biologically meaningful temporal changes directly from paired longitudinal images. It has shown effectiveness across multiple longitudinal imaging applications, including embryo development, wound healing, brain aging, and Alzheimer’s disease. Adapting the LILAC framework, we employed a convolutional neural network (ResNet-18 [10]) with a Siamese architecture [5, 6, 13], which takes a pair of images as input and extracts the corresponding feature representations to predict the correct order of the given longitudinal image pair. For Siamese architecture, ResNet-18 [10] with 3D convolutional layers was used, which demonstrated superior performance compared to simple convolutional neural networks (Supplementary Table A2). The difference between these feature vectors is subsequently fed to a bias-free fully connected layer, enabling the model to learn differences between the two images. The resulting logits are used for binary classification to determine whether the input pair is in the correct or reversed temporal order using the binary cross-entropy loss function.
A4 Training Details
For the temporal-ordering task on image pairs, MR images from fractions were paired in all possible permutations with repetition. Consequently, $n^{2}$ image pairs were generated per patient ($n$: the number of fractions for the patient). In total, this pairing yielded $2,238$ pairs for the training data, $737$ pairs for the validation data, and $732$ pairs for the test data. Ground truth labels were assigned as 1 for the temporally ordered, 0 for the temporally reversed, and 0.5 for the identical pairs.
A curriculum learning approach was adopted as part of the training strategy [4], where a model is gradually exposed to examples in an order from easy to hard to improve learning efficiency and performance. The approach resulted in improved performance compared with training from scratch (Supplementary Table A2).
Models were trained for 100 epochs, with the optimal epoch identified as the one yielding the smallest validation loss. Training and validation losses were monitored to check for convergence and overfitting. Model parameters were optimized using the Adam optimizer [14] to minimize the binary cross-entropy loss between the model outputs and the ground truth. The learning rate was initially set to 0.001 and decayed by a factor of 0.5 every 20 epochs. Training was conducted using the Python PyTorch library on a NVIDIA A40 GPU.
A5 Linear Mixed-Effects model
A linear mixed-effects model was constructed to relate the positive fraction intervals from the first fraction in $F_{1}$-$F_{X}$ ($X>1$) pairs and the corresponding model logits. Because there was no change when comparing the initial time point to itself, intercepts were omitted from the linear mixed-effects model, and only subject-specific random slopes were included. The significance of individual variability was evaluated using a likelihood ratio test comparing the full model with random effects to a reduced model without them.
A6 Grad-CAM
A modified Gradient-weighted Class Activation Mapping (Grad-CAM) [23, 12] was calculated to obtain the saliency maps, which visualize the areas in inputs on which the model largely depends for prediction. Gradients of the model outputs were computed with respect to the final activation maps, and importance was assigned at each spatial location individually. The activation maps were then weighted element-wise by the corresponding gradients to produce the saliency maps. Regions with high Grad-CAM values were analyzed to identify areas likely affected by MR-Linac RT.
For group-level visualization of the saliency maps, $F_{1}$ atlas was constructed from all patients in the test data using ants.registration.build_template module in ANTsPy [26], python implementation of Advanced Normalization Tools (ANTs). The module begins with an initial template, which is the average of all images. All images were nonlinearly registered using SyN (Symmetric Normalization)—ANTs’ diffeomorphic registration model—without rigid or affine alignment, over three iterations. At each iteration, warped images were averaged and blended with the previous template using a blending weight of 0.75 to ensure stable convergence, with deformation updates controlled by a gradient step of 0.2, yielding an unbiased diffeomorphic mean representation of the cohort. Using the constructed atlas, patient-wise saliency maps were transformed into the atlas space, averaged across images, and overlaid for visualization (Figure 4).
A7 Saliency-Restricted-MR Details
Binary masks were generated by thresholding the min-max scaled saliency maps at 0.3. Using these masks, we created bounding boxes that encompassed the thresholded regions. The resulting cropped images were provided to experts for the saliency-restricted-MR evaluation.
A8 Statistical Analysis Details
An independent two-sided t-test (Python scipy.stats.ttest_ind) was conducted for pairwise comparisons of scores (accuracy and AUC) and patient characteristics. The Wilcoxon signed-rank test (Python scipy.stats.wilcoxon) was used to compare organ volume and intensity from $F_{1}$ to $F_{L}$.
Figure A1: Flow diagram of cases, indicating inclusion and exclusion of patients, and random split of the data.
Figure A2: Histogram of days between adjacent fractions.
Figure A3: Pairwise performance of the All-pairs model.(A) Accuracy and (B) AUC of the All-pairs model for different image pairs. Pairs of numbers on the x-axes indicate the fractions at which the paired images were acquired. Cases were grouped by the first fraction and ordered and color-coded by the interval between the paired fractions. Stars indicate statistical significance between compared cases (independent two-sided t-test; ****: $p\leq.0001$).
Figure A4: Peaks of saliency maps of patients in the test data on MR image atlas.
Expert Rationale # of Cases
Areas in the PZ get darker. 85 (56%)
Areas in the TZ get darker. 51 (34%)
TZ features get indistinct. 13 (9%)
Prostate missing but MR signals get darker 2 (1%)
Table A1: Expert rationales for temporal ordering decisions on saliency-restricted-MR images.
Pair Model & Training Scheme ACC AUC
$F_{1}$-$F_{L}$ CNN-3D 0.89 0.96
All CNN-3D, Curriculum learning from the $F_{1}$-$F_{L}$ model 0.89 0.96
CNN-3D, From scratch 0.88 0.95
ResNet-18-3D, From scratch 0.88 0.95
Table A2: Performance of various models and training schemes. CNN-3D consists of four blocks of 3D convolutional layer (number of output channels set as 16, 32, 64, 16, respectively), batch normalization, leaky ReLU, and average pooling. The ResNet-18-3D model trained with curriculum learning, which was used in the study for subsequent analyses, demonstrated superior performance compared with the CNN-3D and the ResNet-18-3D model trained from scratch.